OpenStudy (anonymous):

WILL MEDAL AND FAN How many solutions does the equation have? a+5=1/5(5a+25) a. one solution b. infinite solutions c. no solution

OpenStudy (anonymous):

i think b or c

OpenStudy (anonymous):

2(2x+5)=4(x+3) 4x + 20 = 4x + 12 subtract 4x from both sides 20 = 12 No solution

OpenStudy (anonymous):

idk wich answer it is

OpenStudy (anonymous):

@pizza12

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

no prob:)

OpenStudy (anonymous):

sorry @pizza12 for asking i had no clue i had gotten an answer my computer is lagging right now

OpenStudy (anonymous):

ok it fine but it is c